OpenStudy (anonymous):

What is the value of the expression d – 15 when d = –23?

OpenStudy (some.random.cool.kid):

d - 15 d = -23 -23 - 15 = -38

OpenStudy (some.random.cool.kid):

so the final answer is -38

OpenStudy (anonymous):

Oh okat Thank you some.random.cool.kid